OpenStudy (anonymous):

can someone help me find out what s is? 7=1/6S+11

OpenStudy (anonymous):

where is s in this question?

OpenStudy (anonymous):

after the 6

OpenStudy (anonymous):



OpenStudy (anonymous):

7=1/6S+11 -11 -11 -4=1/6S *1 *1 -4=6S /6 /6 -0.66=S is what I got.

OpenStudy (anonymous):

And love, is also what I got. (song reference)

OpenStudy (anonymous):

@iambatman i love your name (; batmans the is pellet

OpenStudy (anonymous):

the answer was wrong btw it was -24....

OpenStudy (anonymous):

Yup I got - 24.

OpenStudy (anonymous):

@iambatman do you think you could help me on a few more questions?

OpenStudy (anonymous):